The sum of two numbers is twenty-one. Three times the smaller number is two less than twice the larger number. Find the two numbers
Answer by edjones(8007)

Let x be the smaller and y the larger.
a) x+y=21
b) 3x=2y-2
.
a) y=21-x
b)
3x=2(21-x)-2
3x=42-2x-2
5x=40
x=8
y=13
.
Ed